OpenStudy (anonymous):

how do you solve 3u^2-4=0 with no decimal approximations

OpenStudy (amistre64):

by not making it a decimal

OpenStudy (anonymous):

but how do you solve it?

OpenStudy (amistre64):

show me the point where you get stuck at ...

OpenStudy (amistre64):

or walk me thru it to the point your stuck at

OpenStudy (anonymous):

I don't even know where to start

OpenStudy (amistre64):

maybe adding something to both sides would help start it off ....

OpenStudy (amistre64):

or, consider this option ... (what?), minus 4, equals 0 ?

OpenStudy (anonymous):

so would you change it to \[3u^2=4\]

OpenStudy (amistre64):

yes; now consider: 3 times (what?) = 4

OpenStudy (anonymous):

well, wouldn't it have to be a mixed number?

OpenStudy (amistre64):

yes; but lets keep it as an improper fraction

OpenStudy (amistre64):

try dividing both sides by some value ... to get rid of that 3

OpenStudy (anonymous):

ok, so \[u^2=4/3\]

OpenStudy (anonymous):

so you would take the square root of each side?

OpenStudy (amistre64):

very good, now at this point is where people tend to start having troubles ... yes!! but there are 2 results for it: a + and a -

OpenStudy (amistre64):

\[u=\pm\sqrt{4/3}\]

OpenStudy (amistre64):

we can still simplify this more

OpenStudy (anonymous):

yes, the final answer must be in exact form

OpenStudy (amistre64):

most times the do not like to have a sqrt on the bottom of a fraction:\[u=\pm\sqrt{\frac43}\] \[u=\pm\frac{\sqrt4}{\sqrt3}\] \[u=\pm\frac{\sqrt4}{\sqrt3}\frac{\sqrt3}{\sqrt3}\] \[u=\pm\frac{\sqrt4\sqrt3}{\sqrt3^2}\] can you see where its going from there?

OpenStudy (anonymous):

can't you cancel the 3?

OpenStudy (amistre64):

you can cancel the sqrt^2 and leave the 3

OpenStudy (amistre64):

\[u=\pm\frac23 \sqrt3\]

OpenStudy (anonymous):

ohhhh ok!